Mark, Taylor, & Brian's ages are consecutive whole numbers. Mark is the youngest, and Brian is the oldest. The sum of their ages is 111. How old is Taylor?
Answer by TimothyLamb(4379)

z = y + 1
y = x + 1
x + y + z = 111
---
put the system of linear equations into standard form
---
y - z = -1
x - y = -1
x + y + z = 111
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = mark = 36
y = taylor = 37
z = brian = 38
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php